# Scaling Latent Reasoning via Looped Language Models

### v1

@@ 3.4 Stage II: Focused Adaptive Gate Training
Unlike traditional approaches that treat the gating mechanism as an auxiliary component, we directly optimize the exit gate for effective adaptive computation. The core idea of our gate training approach is to teach the model to make termination decisions based on actual performance improvements. We adopt a greedy approach that optimizes the trade-off between improvement via recurrent step and computation efficiency below.
In this stage, the adaptive exit objective must exclusively optimize the gating mechanism without interfering with the language model’s learned representations. We compute the detached loss $\mathcal{L}_{i,\mathrm{stop}}^{(t)}$ at each token position $i$ and define the loss improvement of advancing from step t1t-1 to step $t$ by
$I^{(t)}_{i}=\max(0,\mathcal{L}_{i,\mathrm{stop}}^{(t-1)}-\mathcal{L}_{i,\mathrm{stop}}^{(t)})$ (4)
Intuitively, when $I^{(t)}$ is small, the improvement has stagnated and LoopLM should opt for an early exit. We implement this intuition by computing the ideal continuation probability $w^{(t)}_{i}=\sigma(k\cdot(I^{(t)}_{i}-\tau))$ using $k=50.0$ as the slope and $\tau=0.005$ as the improvement threshold. When $I^{(t)}_{i}>\tau$, $w^{(t)}_{i}$ is large and the model performs the next recurrent step; When $I^{(t)}_{i}\ll\tau$, early exit is favored in step $t$. The adaptive exit loss at step $t$ takes the form of a weighted cross-entropy averaged over the sequence length $M$ between this ideal behavior and the actual gate prediction.
$\mathcal{L}_{\text{adaptive}}^{(t)}=-\frac{1}{M}\sum_{i\leq M}\left[w^{(t)}_{i}\log(1-\lambda^{(t)}_{i})+(1-w^{(t)}_{i})\log(\lambda^{(t)}_{i})\right$] (5)
The total adaptive training loss is averaged across all recurrent steps:
@@
The adaptive loss (5) trains the gate in step $t$ by matching its weights $\lambda^{(t-1)}_{i}$ with the ideal probability $w^{(t)}_{i}$. This formulation penalizes two failure modes simultaneously:
  • Underthinking, where the gate wants to stop when it should continue (when the ideal continuation probability $w^{(t)}_{i}$ is large, but the early stop gate $\lambda^{(t-1)}_{i}$ is also large);
  • Overthinking, where the gate wants to continue when it should stop (when the ideal continuation probability $w^{(t)}_{i}$ is small, but $\lambda^{(t)}_{i}$ is also small).
By optimizing (5), the gates learn to greedily search for the optimal exit step, to tradeoff computation efficiency for better performance. For empirical evaluations, see Section˜5.4.1.
@@ 4 Training Looped Language Models
Figure 4: The Ouro model training pipeline. The process starts with a common Warmup and an initial 3T token Stable Training phase. The model is then split into two streams: one ‘Keep 1.3B’ (resulting in Ouro-1.4B) and one ‘Upcycle 2.6B’ (resulting in Ouro-2.6B). Both streams independently undergo an identical subsequent four-stage training process: a second Stable Training (3T tokens), CT Annealing (1.4T tokens), LongCT (20B tokens), and Mid-Training (300B tokens). This 7.7T token pre-training pipeline produces the base models (Ouro-1.4B and Ouro-2.6B), which are then passed through a final Reasoning SFT stage to create the Ouro-Thinking models.
Our training pipeline for the Ouro model family is a multi-stage process, as illustrated in Figure 4. The process begins with a common warmup stage, followed by an initial Stable Training phase on 3T tokens. After this, the model is split into two variants: a 1.4B parameter model and a 2.6B model created via upcycling. Both variants then undergo an identical series of four subsequent training stages: a second Stable Training phase (3T tokens), CT Annealing (1.4T tokens), LongCT (20B tokens), and Mid-Training (300B tokens). This comprehensive pipeline, totaling 7.7T tokens of training data, produces our base models, Ouro-1.4B and Ouro-2.6B. Finally, these base models are further refined through a specialized Reasoning SFT (Supervised Fine-Tuning) stage to create the final, reasoning-focused models: Ouro-1.4B-Thinking and Ouro-2.6B-Thinking. This section details the model architecture, data composition, and specific configurations used in each of these training stages.
4.1 Transformer Architecture
Our Ouro models are built upon the standard decoder-only Transformer architecture [31], prioritizing a clean implementation of the looped computation mechanism without extraneous modifications. The core architecture consists of a stack of identical Transformer blocks, which are applied recurrently.
Each block uses Multi-Head Attention (MHA) with Rotary Position Embeddings (RoPE) [32] to handle sequence order. For computational efficiency, The feed-forward network (FFN) in each block utilizes a SwiGLU activation [33]. To enhance training stability, which is especially critical for deep recurrent computation, we employ a sandwich normalization structure. This places an RMSNorm layer before both the attention and FFN sub-layers, an approach noted in prior literature to improve stability in loop transformers [15].
Ouro 1.4B 1.4B 24 2048 MHA SwiGLU RoPE 49,152
Ouro 2.6B 2.6B 48 2048 MHA SwiGLU RoPE 49,152
For both models, we use a shared vocabulary of 49,152 tokens, reused from the SmolM2 [34] model. This tokenizer is optimized for languages with a Latin alphabet and code and contains no Chinese tokens. Our inclusion of Chinese data in Stage 1 thus resulted in highly inefficient tokenization and poor performance. Consequently, we removed all Chinese data from Stage 2 onwards to focus our training budget on English and code. This limited vocabulary may also impose constraints on the model’s advanced mathematical reasoning capabilities due to a potential lack of specialized symbols. This shift is reflected in the data compositions detailed in the following sections.
4.2 Data
As data defines the capability boundaries of large foundational models, our model is trained on a diverse collection of datasets spanning multiple domains and stages, including web data, mathematical content, code, and long-context documents, enabling it to perform acquire fundamental language understanding and perform advanced reasoning, coding, and long-context understanding through a unified training pipeline. In addition to standard web crawl datasets, we adopt specialized datasets for mathematical reasoning and code generation to further enhance the model’s capabilities for complex problem-solving. In Table˜2, we summarize the composition and quantity of our training data across different stages. In the following sections, we detail our dataset sources, preparation protocols, and data mixing strategies.
Data Source Stage # Tokens (B) # Used Tokens (B)
Nemotron-CC (Web Data) Stage 1 6386 4404
MAP-CC (Web Data) Stage 1 800 780
Ultra-FineWeb-zh (Web Data) Stage 1 120 120
OpenCoder-pretrain Stage 1 450 450
MegaMath-web Stage 1 247 246
MegaMath-high-quailty Stage 2 64 64
Nemotron-$C$-Math-v1 Stage 2 210 210
Nemotron-Code Stage 2 53 53
Nemotron-SFT-Code Stage 2 48 48
Nemotron-SFT-General Stage 2 87 87
OpenCoder-Annealing Stage 2 7 7
ProLong-64K Stage 3 20 20
Mid-training SFT Mix Stage 4 182 90
Table 3: Data composition for Stage 1 (Pre-training Phase I & II). Total dataset size: 6T tokens.
Proportion (%) 73.4 13.0 2.0 7.5 4.1
4.2.1 Data Composition
The capabilities of modern language models primarily stem from their training data, and this principle holds true for our model as well. To ensure reproducibility, our training corpus is entirely composed of open-source datasets, with data statistics summarized in Table 3. We partition the data into four distinct subsets, called stages, each employing different data construction strategies that align with the Warmup-Stable-Decay (WSD) [35] learning rate scheduler, which is widely adopted in modern language model pretraining.
Stage 1: Pre-training
The pre-training stage supports the warmup and stable phases of training. The dataset is primarily composed of Web CommonCrawl (CC) data. Since our objective is to train the model on more than 2T tokens, many commonly used open-source datasets would not suffice: Fineweb-Edu [36] provides 1.3T tokens, and DCLM [37] offers 2.6T tokens. We select Nemotron-CC [38] (6.3T tokens) as the main dataset for the stable phase due to its large scale and suitability for our training requirements. To provide the model with basic Chinese proficiency, we include Ultra-FineWeb-zh [39] and MAP-CC [40]. Additionally, to enhance coding and mathematical abilities, we incorporate OpenCoder [41] and MegaMath [42].
Stage 2: Continual Training (CT) Annealing
The CT annealing stage incorporates higher-quality data to enhance the model under the annealing learning rate. We construct our dataset using the high-quality subset of Nemotron-CC as the base. To further strengthen mathematical, coding, and general capabilities, we incorporate the high-quality subset of MegaMath, Nemotron-CC-Math-v1 [43, 44], OpenCoder-Annealing [41], Nemotron-Pretraining-Code-v1 [44], and Nemotron-Pretraining-SFT-v1 [44].
Data Source Proportion (%)
Nemotron-$C$-high-quailty 66.5
Nemotron-$C$-Math-v1 15.0
MegaMath-high-quailty 4.6
OpenCoder-LLM/opc-annealing-corpus 0.5
Nemotron-Pretraining-Code-v1/Synthetic-Code 3.8
Nemotron-Pretraining-SFT-v1/Nemotron-SFT-Code 3.4
Nemotron-Pretraining-SFT-v1/Nemotron-SFT-General 6.2
Stage 3: Long Context Training (LongCT)
Stage 4: Mid-training
[…]
Table 5: Training recipe for both Ouro 1.4B and 2.6B.
@@ 4.2.1 Data Composition
4.3 Pre and Mid-Training
We adopt a multi-stage training strategy using a dynamic mixture of the curated data described, specifically: a Pre-training stage (split into two phases with different recurrent configurations), a CT Annealing stage for quality enhancement, a LongCT stage for context extension, and a Mid-training stage for advanced capability refinement. Throughout our training pipeline, we train the model with a maximum of 4 recurrent steps.
4.3.1 Training Stability and Adaptive Configuration
During training, we prioritized stability over aggressive scaling, making several key adjustments based on empirical observations of training dynamics. These decisions were critical for achieving stable convergence with recurrent architectures, which exhibit different optimization characteristics compared to standard transformers.
@@ 4.3.1 Training Stability and Adaptive Configuration
Our initial experiments with 8 recurrent steps in Stage 1a reveal training instabilities, including loss spikes and gradient oscillations. We hypothesize this stems from the compounding gradient flow through multiple recurrent iterations, which can amplify small perturbations. Consequently, we reduced the recurrent steps from 8 to 4 in Stage 1b, finding this sweet spot balanced computational depth with training stability.
@@
We strategically reduced $\beta$ from 0.1 in Stage 1a to 0.05 in subsequent stages. This reduction serves dual purposes: (1) it decreases the conflicting gradients between task loss and the KL penalty, leading to more stable optimization, and (2) it reduces the “pull” from the uniform prior, allowing the model greater freedom to explore beneficial depth patterns without being artificially constrained. This adjustment was crucial for maintaining stable training dynamics while enabling the model to learn effective depth allocation.
4.3.2 Stage-wise Training Details
We utilize the flame [46] framework for performing pretraining, built upon torchtitan [47]. To fully utilize our resources, we adopt an upcycling strategy that enables efficient scaling of model capacity during training.
  • Stage 1a: Pre-training Phase I (Exploration Phase). We initially train the model on 3T tokens of web data from Nemotron-CC with 8 recurrent steps. The training uses the Warmup-Stable-Decay (WSD) learning rate scheduler with a peak learning rate of $3\times 10^{-4}$. The sequence length is set to 4K tokens with an initial batch size of 4M tokens, gradually increased to 8M for stability. During this phase, we observed training instabilities that prompted our subsequent architectural adjustments.
  • Stage 1b: Pre-training Phase II with Stability-Driven Upcycling. After identifying stability issues in Stage 1a, we implemented an architectural pivot: reducing recurrent steps from 8 to 4. To maintain computational efficiency while improving stability, we split our approach into two variants:
    • Variant 1: A 1.4B parameter model maintaining 24 layers with 4 recurrent steps
    • Variant 2: A 2.6B parameter model created through layer stacking (48 layers) with 4 recurrent steps
    The recurrent nature of our architecture makes this upcycling process particularly smooth, as the shared weights across iterations naturally facilitate layer duplication without the typical instabilities seen in standard transformer upcycling. Both variants are trained on an additional 3T tokens with the stabilized configuration. The data composition is carefully balanced as shown in Table 3.
  • Stage 2: CT Annealing. Building on the stable foundation from Stage 1b, we enhance the model with higher-quality data while annealing the learning rate to $3\times 10^{-5}$. The training corpus comprises 1.4T tokens with increased emphasis on mathematical and coding capabilities. >We extend the sequence length to 16K tokens, exceeding the length of most samples to minimize truncation and better utilize the enhanced data quality. The recurrent steps remain at 4, having proven optimal for the stability-performance trade-off. The data composition is carefully balanced as shown in Table 4.
  • Stage 3: LongCT. This stage focuses on extending the model’s context window capabilities. We train on 20B tokens from the ProLong-64K dataset with sequences of 64K tokens, maintaining the batch size at 8M tokens. The reduced KL coefficient ($\beta=0.05$) continues to provide stable training dynamics even with these extended sequences.
  • Stage 4: Mid-training. The final stage leverages 90B tokens of extremely high-quality SFT data, consisting of both <Question, Answer> and <Question, CoT, Answer> samples. All SFT-style data is converted to ChatML format to facilitate subsequent post-training alignment. The learning rate is further reduced to $1\times 10^{-5}$ with a cosine scheduler to help the model better absorb on this diverse, high-quality dataset.
@@ 4.3.2 Stage-wise Training Details
We empirically found that recurrent architectures require smaller learning rates compared to standard transformers of equivalent parameter count. While resource constraints prevented exhaustive hyperparameter search, our chosen rates represent conservative values that prioritized stable convergence over potentially faster but riskier optimization trajectories.
@@
The sequence length is progressively increased across stages: 4K tokens for both pre-training phases, 16K for continual training with learning rate annealing, 64K for long-context training, and 32K for mid-training. This gradual progression helps maintain stability while enhancing the training throughput and expanding the model’s long-context capability.

### v2

@@ 3.4 Stage II: Focused Adaptive Gate Training
In this stage, we freeze the LM parameters and train only the exit gate to make termination decisions based on realized performance gains. We use a greedy signal that balances marginal improvement from an extra loop against additional compute.
To ensure the gate does not alter LM representations, we compute a detached per-step loss $\mathcal{L}_{i,\mathrm{stop}}^{(t)}$ at each token $i$ and define the loss improvement from step t1t\!-\!1 to $t$ as
$I^{(t)}_{i}=\max\!\big(0,\ \mathcal{L}_{i,\mathrm{stop}}^{(t-1)}-\mathcal{L}_{i,\mathrm{stop}}^{(t)})$ (5)
where larger $I_{i}^{(t)}$ indicates ongoing improvement; a smaller value indicates that gains have stalled and LoopLM should opt for an early exit. We implement this by computing the ideal continuation probability, a training label that indicates whether to continue (near 1) or exit (near 0):
$w^{(t)}_{i}=\sigma(k\cdot(I^{(t)}_{i}-\gamma))$
with slope $k=50.0$ and threshold $\gamma=0.005$, so that $w^{(t)}_{i}\!\approx\!1$ recommends continuing and $w^{(t)}_{i}\!\approx\!0$ recommends exiting the loop. The adaptive exit loss at step $t$ takes the binary cross-entropy between the gate’s predicted continuation probability $1-\lambda^{(t)}_{i}$ and the ideal label $w^{(t)}_{i}$, averaged over the sequence length $M$:
$\mathcal{L}^{(t)}_{\text{adaptive}}=-\frac{1}{M}\sum_{i=1}^{M}\!\Big[w^{(t)}_{i}\,\log\bigl(\underbrace{1-\lambda^{(t)}_{i}}_{\begin{subarray}{c}\text{predicted}\\ \text{continuation}\end{subarray}}\bigr)+\bigl(1-w^{(t)}_{i}\bigr)\,\log\bigl(\underbrace{\lambda^{(t)}_{i}}_{\begin{subarray}{c}\text{predicted}\\ \text{exit}\end{subarray}}\bigr)\Big].$ (6)
The total adaptive loss averages across recurrent steps:
@@
The adaptive loss in Equation˜6 trains the gate at step $t$ to match its predictions to the ideal behavior derived from actual performance improvements:
  • Predicted probabilities: The gate generates $\lambda^{(t)}_{i}$ (exit probability) and $1-\lambda^{(t)}_{i}$ (continuation probability)
  • Target labels: The ideal behavior is encoded as $w^{(t)}_{i}$ (target continuation probability) and $1-w^{(t)}_{i}$ (target exit probability)
This formulation penalizes two failure modes simultaneously:
[…]
Optimizing Equation˜6 trains the gate to choose a greedy exit step that trades additional compute for measured improvement. For empirical evaluations, see Section˜5.4.1.
@@ 4 Training Looped Language Models
Figure 4: End-to-end Ouro training pipeline: shared warmup → Stable Training → forks into a 1.4B retained path and a 2.6B upcycled path → four shared stages → Reasoning SFT to produce Ouro-Thinking.
Our end-to-end training pipeline for the Ouro model family is shown in Figure 4. In total 7.7T tokens are used to train the base models Ouro-1.4B and Ouro-2.6B. A final Reasoning SFT (Supervised Fine-Tuning) yields the Ouro-1.4B-Thinking and Ouro-2.6B-Thinking variants. This section details the architecture, data composition, and specific configurations used in each of these training stages. A high-level overview of the training recipe of the first four stages is given in Table˜1.
Table 1: Training recipe for Ouro 1.4B and 2.6B.
@@
4.1 Transformer Architecture
The Ouro models use a standard decoder-only Transformer [33], prioritizing a clean implementation of the looped computation mechanism without extraneous modifications. The core architecture consists of a stack of Transformer blocks applied recurrently. Each block uses Multi-Head Attention (MHA) with Rotary Position Embeddings (RoPE) [34]. The feed-forward network (FFN) in each block utilizes a SwiGLU activation [35]. To enhance training stability, which is especially critical for deep recurrent computation, we employ a sandwich normalization structure, placing an RMSNorm layer before both the attention and FFN sub-layers [17]. Both models use a 49,152-token vocabulary from the SmolLM2 model [36]. This tokenizer is optimized for code and Latin-alphabet language. Architectural details are summarized in Table˜2.
Ouro 1.4B 1.4B 24 2048 MHA SwiGLU RoPE 49,152
Ouro 2.6B 2.6B 48 2048 MHA SwiGLU RoPE 49,152
4.2 Data
Data sets the capability bounds of foundation models. Our corpus spans web text, mathematics, code, and long-context documents across multiple stages, building core language understanding while strengthening reasoning, coding, and long-context skills. Beyond standard web crawls, we include targeted datasets for mathematical reasoning and code generation to improve complex problem solving. Table˜3 summarizes composition and scale at each training stage.
Data Source Stage # Tokens (B) # Used Tokens (B)
Nemotron-CC (Web Data) Stage 1 6386 4404
MAP-CC (Web Data) Stage 1 800 780
Ultra-FineWeb-zh (Web Data) Stage 1 120 120
OpenCoder-pretrain Stage 1 450 450
MegaMath-web Stage 1 247 246
MegaMath-high-quailty Stage 2 64 64
Nemotron-$C$-Math-v1 Stage 2 210 210
Nemotron-Code Stage 2 53 53
Nemotron-SFT-Code Stage 2 48 48
Nemotron-SFT-General Stage 2 87 87
OpenCoder-Annealing Stage 2 7 7
ProLong-64K Stage 3 20 20
Mid-training SFT Mix Stage 4 182 90
Table 4: Data composition for Stage 1 (Stable Training I & II). Total dataset size: 6T tokens.
Proportion (%) 73.4 13.0 2.0 7.5 4.1
To ensure reproducibility, our training corpus is composed entirely of open-source datasets, with data statistics summarized in Table 4. We partition the data into four stages, each with construction strategies aligned to the Warmup-Stable-Decay (WSD) learning rate scheduler [37] commonly used in modern pre-training.
Stage 1: Pre-training
This stage supports the warmup and stable phases of training. The corpus is primarily composed of Web CommonCrawl (CC) data. Because we sought to train the model on >2T tokens, many popular open corpora are too small (e.g., Fineweb-Edu at 1.3T tokens [38], DCLM at 2.6T tokens [39]). We therefore use Nemotron-CC [40] (6.3T tokens) as the main dataset for the stable-phase. To provide the model with basic Chinese proficiency, we include Ultra-FineWeb-zh [41] and MAP-CC [42]. However, without Chinese vocabulary in the tokenizer, characters would be fragmented into multiple byte-level sub-tokens, so we removed Chinese from Stage 2 onwards. To enhance coding and mathematical abilities, we incorporate OpenCoder [43] and MegaMath [44]. See Table˜4 for dataset proportions in further detail.
Stage 2: Continual Training (CT) Annealing
The CT annealing stage incorporates higher-quality data to enhance the model under the annealing learning rate. Token sequence length is extended to 16K tokens, exceeding the length of most samples to minimize truncation. We construct the corpus from the high-quality subset of Nemotron-CC and augment with HQ MegaMath, Nemotron-CC-Math-v1 [45, 46], OpenCoder-Annealing [43], Nemotron-pre-training-Code-v1 [46], and Nemotron-pre-training-SFT-v1 [46]. Data composition is provided inTable˜5.
Data Source Proportion (%)
Nemotron-$C$-high-quailty 66.5
Nemotron-$C$-Math-v1 15.0
MegaMath-high-quailty 4.6
OpenCoder-LLM/opc-annealing-corpus 0.5
Nemotron-pre-training-Code-v1/Synthetic-Code 3.8
Nemotron-pre-training-SFT-v1/Nemotron-SFT-Code 3.4
Nemotron-pre-training-SFT-v1/Nemotron-SFT-General 6.2
Stage 3: Long Context Training (LongCT)
Stage 4: Mid-training
[…]
4.3 Training Stability and Adaptive Configuration
We use the flame [48] framework for pre-training, built on torchtitan [49]. During training, we prioritized stability over aggressive scaling, making several key adjustments based on empirical observations of training dynamics. These decisions were critical for achieving stable convergence with recurrent architectures, which exhibit different optimization characteristics compared to standard transformers.
@@ 4.3 Training Stability and Adaptive Configuration
Our initial experiments with 8 recurrent steps in Stage 1a (Stable Training I) led to loss spikes and gradient oscillations. We hypothesize this stems from compounded gradient flow through multiple recurrent iterations, which can amplify small perturbations. Consequently, we reduced the recurrent steps from 8 to 4 in Stage 1b (Stable Training II in Figure˜4), which balanced computational depth with training stability.
@@
We strategically reduced $\beta$ in Equation˜4 from 0.1 in Stage 1a to 0.05 in later stages. This reduction serves dual purposes: (1) it decreases the conflicting gradients between task loss and the KL penalty, leading to more stable optimization, and (2) it reduces the “pull” from the uniform prior, allowing the model greater freedom to explore beneficial depth patterns without being artificially constrained. This adjustment allowed the model to learn useful depth patterns without undue constraint.
@@
We empirically found that recurrent architectures require smaller learning rates than parameter-matched Transformers. Given compute constraints, we did not run exhaustive LR sweeps, but instead, adopted conservative rates that prioritized stable convergence over potentially faster but riskier schedules.
@@
The sequence length is progressively increased across stages: 4K tokens for both pre-training phases, 16K for CT annealing, 64K for long-context training, and 32K for mid-training. This progression stabilizes optimization while expanding context capacity with training throughput.
4.3.1 Stage-wise Training Details
  • Stage 1a: Pre-training Phase I (Exploration). We initialize training with 8 recurrent steps. The learning rate follows a Warmup-Stable schedule with a peak of $3\times 10^{-4}$. The sequence length is 4K tokens with an initial batch size of 4M tokens, gradually increased to 8M for stability. During this phase, we observed training instabilities that prompted subsequent architectural adjustments.
  • Stage 1b: Pre-training Phase II with Stability-Driven Upcycling. After identifying stability issues in Stage 1a, we reduced the recurrent steps from 8 to 4. To maintain computational efficiency while improving stability, we split our approach into two variants:
    • 1.4B Ouro: Uses the original 24 pre-trained layers with 4 recurrent steps
    • 2.6B Ouro: Upcycles 24 layers to 48 via layer duplication with 4 recurrent steps
    The recurrent nature of our architecture makes this upcycling process particularly smooth, as the shared weights across iterations naturally facilitates layer duplication without the typical instabilities seen in standard transformer upcycling.
  • Stage 2: CT Annealing. The learning rate is annealed to $3\times 10^{-5}$ while exposing the model to high-quality training data. The recurrent steps remain at 4, having proven optimal for the stability-performance trade-off. The data composition is carefully balanced as shown in Table 5.
  • Stage 3: LongCT. The batch size is held at 8M tokens. The reduced KL coefficient ($\beta=0.05$) continues to provide stable training dynamics even with the 64K-length sequences.
  • Stage 4: Mid-training. The learning rate is further reduced to $1\times 10^{-5}$ with a cosine scheduler to help the model better absorb on this diverse, high-quality dataset.